OpenStudy (anonymous):

8. The probability of the complement of an event is ____ less than the probability of the event itself. A. sometimes B. always C. never

OpenStudy (anonymous):

please someone help

jimthompson5910 (jim_thompson5910):

let's say event A has a 20% chance of happening in other words, P(A) = 0.2 the complement of event A is the opposite of it (ie the event of A not happening) So the probability of this is 1 - P(A) = 1 - 0.2 = 0.8 So we can see that the complement has a better chance of happening (higher probability)

jimthompson5910 (jim_thompson5910):

we can easily swap the probabilities for P(A) and P(not A) so P(A) = 0.8 for instance and P(not A) = 0.2

jimthompson5910 (jim_thompson5910):

there's no rule that says that the complement event has to have a larger or smaller probability